OpenStudy (anonymous):

Transformations help- pre calc. Hi I got a little confused with vertical stretches when I was asked this question: "Replace y with 2y in the function y= f(x)" The answer is a vertical stretch by a factor of 1/2. I know with horizontal stretches you use the reciprocal. But with this question instead of 'a' were given y? o.o. So 2= f(x)? and then I'm not sure where to go from there

OpenStudy (lastdaywork):

Can you state an example? It's hard to comprehend what you wanna know.

OpenStudy (anonymous):

oh uhmmm, I'll write the full question. and some other examples: Describe what happens to the graph of a function y= f(x0 after the following changes are made to its equation: A) replace y with 2y <--- my question Example: B) Replace x with 1/4x = horizontally stretched by a factor of 4

OpenStudy (anonymous):

f(x) *

OpenStudy (lastdaywork):

"A) replace y with 2y <--- my question" Each y-coordinate of original graph gets divided by 2. In other words "a vertical stretch by a factor of 1/2"

OpenStudy (lastdaywork):

"B) Replace x with (1/4)x = horizontally stretched by a factor of 4" That's correct

OpenStudy (anonymous):

If it was written in an equation form how would look like? Do I leave the y isolated or? o.o or is "y= 2y" actually the f in y=f(x)? Sorry I'm pretty bad at math :P

OpenStudy (anonymous):

I mean it would be y= af(x)? o.o

OpenStudy (lastdaywork):

Please talk in terms of examples; it's hard to understand your words (no kidding) "y= af(x)" You can re-write it as (1/a)y = f(x)

OpenStudy (anonymous):

ohhh that's all I needed to know. Thanks :P

OpenStudy (lastdaywork):

:)